OpenStudy (rock_mit182):

A manufacturing company that has just located in a small community will pay two million dollars per year in salaries. It has been estimated that 60% of these salaries will be spent in the local area, and 60% of the money spent will again change hands within the community. This process, called the multiplier effect, will be repeated ad infinitum. Find the total amount of local spending that will be generated by company salaries.

OpenStudy (whpalmer4):

Do you know the formula for the sum of a geometric sequence? What is the limit as the number of terms goes to infinity?

OpenStudy (rock_mit182):

no, could you show me

OpenStudy (whpalmer4):

The sum of a geometric sequence is \[\sum _{k=0}^{n-1} a r^k = \frac{a \left(1-r^n\right)}{1-r}\] What happens to that as \(n\rightarrow\infty\)? (Remember that \(|r| < 1\) if this converges!)

OpenStudy (rock_mit182):

It seems like converges cause the numerator becomes negative, and each tiem more larger

OpenStudy (rock_mit182):

oh man but I'm not good at this at all. .. so what is the next step ?

OpenStudy (rock_mit182):

in fact i don't know what is a and r here ( in the problem)

OpenStudy (whpalmer4):

As \(n\rightarrow\infty\), that \(r^n\) term goes to 0, leaving us with just \(\large a*\frac{1}{1-r}\). \(a\) is the amount of money invested...

OpenStudy (rock_mit182):

so you mean\[\lim_{n \rightarrow \infty} r ^{n} = 0\] cause r ^ -10 ^6 and so on is more near to zero

OpenStudy (rock_mit182):

ok understood, but how is this related with the problem ?

OpenStudy (rock_mit182):

I guess now i have to plug n chuff values on the equation left

OpenStudy (rock_mit182):

\[\sum_{0}^{\infty}=\frac{ 60 }{ 1-60 }\] right ?

OpenStudy (whpalmer4):

sorry, juggling too many questions at once! In a geometric sequence, we have terms \({a, a*r, a*r*r, a*r*r*r, ... a*r^n}\) For our application, \(a = \$2,000,000\) and \(r = 60\% = 0.60\) So the first round of circulation, $2 million is spent. The second round, $2 million *0.6 = $1.2 million is spent. The third round, $1.2 million * 0.6 = $0.72 million, and so on.

OpenStudy (whpalmer4):

We can factor out that value of \(a\): \[\sum _{k=0}^{n-1} a r^k = a \sum_{k=0}^{n-1} r^k = a\frac{ \left(1-r^n\right)}{1-r}\] As \(n\rightarrow\infty\), the fraction approaches \(\large\frac{1}{1-r}\) so the total money spent and spent and spent is just \[a\frac{1}{1-r} = \$2000000(\frac{1}{1-0.60}) = \] You can test this for reasonability at the other end by saying that none of the money is respent by setting \(r = 0\). If you do that, what do you get?

OpenStudy (rock_mit182):

oh i see it behaves like a linear operator

OpenStudy (whpalmer4):

I'll have to take your word for that :-) Here's a plot of the effect as \(r\) goes from \(0\rightarrow 1\) Clearly, if you can boost that percentage of money recirculating in the local community even just a little bit, it has a bit impact!

OpenStudy (whpalmer4):

big, not bit

OpenStudy (rock_mit182):

awesome you know a lot of math

OpenStudy (rock_mit182):

Thanks bro

OpenStudy (whpalmer4):

You're welcome!